### v1

# From Seeing to Doing: Bridging Reasoning and Decision for Robotic Manipulation

Abstract
Achieving generalization in robotic manipulation remains a critical challenge, particularly for unseen scenarios and novel tasks. Current Vision-Language-Action (VLA) models, while building on top of general Vision-Language Models (VLMs), still fall short of achieving robust zero-shot performance due to the scarcity and heterogeneity prevalent in embodied datasets. To address these limitations, we propose FSD (From Seeing to Doing), a novel vision-language model that generates intermediate representations through spatial relationship reasoning, providing fine-grained guidance for robotic manipulation. Our approach combines a hierarchical data pipeline for training with a self-consistency mechanism that aligns spatial coordinates with visual signals. Through extensive experiments, we comprehensively validated FSD’s capabilities in both “seeing” and “doing”, achieving outstanding performance across 8 benchmarks for general spatial reasoning and embodied reference abilities, as well as on our proposed more challenging benchmark VABench. We also verified zero-shot capabilities in robot manipulation, demonstrating significant performance improvements over baseline methods in both SimplerEnv and real robot settings. Experimental results show that FSD achieves 54.1% success rate in SimplerEnv and 72% success rate across 8 real-world tasks, outperforming the strongest baseline by 30%. More visualizations are available on website.
1 Introduction
A driving force behind robotics research is the pursuit of generalization: creating agents capable of versatile action across diverse robotic platforms, extending beyond familiar tasks, objects, and environments while adapting to dynamic visual inputs. Current approaches [21, 3] leverage pre-trained Vision-Language Models (VLMs) and transform them into Vision-Language-Action Models (VLAs) using large-scale embodied datasets. This enables systems to interpret natural language instructions and generate robotic manipulation actions [2]. The intention is to capitalize on the generalization capabilities of VLMs pre-trained on internet-scale data, with the hope that resulting VLAs will adapt to novel scenarios involving unseen objects and tasks. However, empirical evidence [72, 67, 35] demonstrates that directly applying the generalization power of VLMs falls short of achieving strong zero-shot performance on completely novel tasks.
Figure 1: Overview of FSD. FSD unlocks visual aids reasoning and generation through Spatial Relationship-Focused CoT, demonstrating exceptional generalization capabilities that enable zero-shot robot manipulation and achieving remarkable performance across multiple benchmarks.
We attribute the limited generalization in VLA-based systems to scarcity and heterogeneity of datasets. Robotics data remains limited compared to language and vision datasets, preventing similar scaling laws [18, 29]. Despite growth in embodied datasets [40], their insufficient coverage and diversity prevent robust zero-shot generalization. Additionally, robotic embodiment heterogeneity [54] causes significant variation in VLA outputs, making end-to-end supervised learning from vision and language to diverse action outputs a potentially unrealistic path toward generalization.
We introduce a novel pipeline addressing generalization challenges in robotic manipulation. Our approach leverages VLMs’ visual understanding capabilities, augmented with step-by-step reasoning to extract unified mid-level structural representations independent of robot embodiment. This representation is key to generalizing learning across diverse physical interactions and dynamic behaviors. Specifically, this mid-level representation includes spatial affordance boxes/points and visual traces, each represented as marked coordinates within visual images. These visual aids provide expressive yet compact spatial information that enables effective reasoning and decisions, overcoming both scarcity and heterogeneity limitations. We introduce FSD (From Seeing to Doing), a model that generates visual intermediate representations through spatial understanding and reasoning (Fig. 1). The FSD model comprises three key components: (1) Spatial Relationship-Focused Visual Chain-of-Thought (SrCoT), which conducts multi-step reasoning anchored by object coordinates and spatial relationships, treating visual aid generation as a reasoning process; (2) A hierarchical data pipeline combining large-scale embodied datasets with common sense data, establishing a weak-to-strong capability enhancement training process; and (3) A self-consistency mechanism that aligns understanding and generation by binding spatial coordinates with specific visual signals. We employ supervised fine-tuning of VLMs using annotated spatial COT datasets for visual aids generation, then action executors follow spatial affordances and visual traces through simple planning methods to complete action execution.
Our FSD model generalizes effectively to new instructions, and scenes through enhanced reasoning abilities. Our contributions include: (1) A novel paradigm bridging VLM reasoning with embodied decisions via visual aids; (2) The SrCoT method enabling multi-step reasoning for visual aid generation and guiding zero-shot manipulation; (3) Our weak-to-strong spatial reasoning and visual aids datasets, along with VABench, a manually annotated challenging benchmark for visual aids generation; (4) Superior performance across 8 benchmarks in spatial reasoning, free space reference, and visual aids generation, with zero-shot deployment achieving 54.1% success in SimplerEnv and 73% in 8 real-world tasks, outperforming mainstream VLA models by 30%.
2 Related Work
Spatial Understanding and Reasoning with VLMs Spatial understanding and reasoning [32, 48, 8, 28, 44] require VLMs to infer spatial information beyond 2D RGB images, a capability crucial for embodied AI applications such as navigation [48, 12, 22] and manipulation [66]. Recent advances include SpatialVLM [5], which explicitly incorporates spatial primitives and coordinate systems to enhance geometric reasoning capabilities. Similarly, SpatialRGPT [7] and SpatialBot [4] improve spatial capabilities through more precise spatial relationship modeling. FSD specifically targets embodied manipulation scenarios by enhancing spatial reasoning capabilities through novel SrCoT mechanisms and sophisticated self-consistency alignment techniques.
Visual Chain-of-thought Reasoning The integration of Chain-of-thought (CoT) [56] and its variants [69, 63, 62] has significantly enhanced LLM reasoning abilities through structured step-by-step processes. For the multimodal reasoning challenge, researchers have developed various CoT approaches [37, 70, 61, 58] that establish appropriate reasoning anchors and extend reasoning pathways. Shikra [6] improves referential expression by incorporating specific visual regions into the reasoning process, while VoCoT [27] extends visually-grounded reasoning chains. VisualCoT [46] provides a comprehensive benchmark validating CoT effectiveness in multi-hop image reasoning tasks. EmbodiedCoT [67] pioneered CoT application in embodied AI by enhancing intermediate reasoning through fine-tuning OpenVLA [21]. In contrast, FSD uniquely integrates visual-spatial reasoning using spatial relationship graphs as reasoning anchors.
Visual Aids Empowered Robotic Manipulation Extracting embodiment-agnostic visual aids to enhance training efficiency has emerged as a promising paradigm in robotic manipulation. Numerous studies [1, 57, 59, 72, 65] have explored robotic policy learning based on visual traces, though most remain confined to specific tasks with cross-embodiment applicability. LLaRVA [39] advances this field by predicting visual traces to better align visual and action spaces for robot learning, compiling an impressive large visual trace dataset, yet struggles to generalize to novel downstream tasks without task-specific fine-tuning. Spatial affordance represents another effective visual aid, with several works [66, 38, 42, 48, 16, 60, 26] demonstrating its utility in robotic manipulation tasks. Our key insight is that the scarcity of embodied data fundamentally limits purely data-driven visual aid prediction approaches, impeding zero-shot generalization to unseen scenarios. Therefore, we employ a reasoning-driven paradigm to activate general spatial abilities and enhance generalization.
3 Bridging Reasoning and Decision through Visual Aids Generation
To harness VLMs’ visual perception capabilities for cross-domain generalization, we introduce FSD (From Seeing to Doing)—a model exhibiting robust spatial reasoning and generation abilities. We first establish visual aids through spatial affordances and visual traces, then develop Spatial Relationship-Focused CoT (SrCoT), which leverages object-centric coordinates and their spatial relationships as reasoning anchors. Supporting this approach requires precise spatial understanding and complex instruction-following capabilities to generate coordinate representations. We implement a progressive weak-to-strong data construction pipeline across five capability levels, complemented by a self-consistency alignment mechanism that enhances understanding and generation abilities.
3.1 Definition of Visual Aids
Figure 2: Diagrams of Visual Aid Types
As shown in Fig. 2, FSD utilizes three visual aids of increasing complexity, all defined within normalized image coordinates $\mathbf{x}=(p,q)\in[0,1000]^{2}\subset\mathbb{R}^{2}$: Spatial affordance boxes $\mathcal{B}=[x_{1},y_{1},x_{2},y_{2}]$ define target regions for object placement. For instructions like "Put the sushi into the silver pot," the model must infer coordinates for unmarked free space—beyond standard detector capabilities. Spatial affordance points $\mathcal{P}=\{(x_{i},y_{i})\mid i=1,2,...,n\}$ provide more precise and flexible placement with reduced redundancy. Object-centric visual traces $\boldsymbol{\tau}=\{\mathbf{x}_{t}\mid t=1,2,...,T\}$ represent ordered coordinate sequences that describe manipulation trajectories, where $T$ denotes the sequence length. These traces enable complex instruction execution, cross-embodiment transfer, and collision avoidance. We implement these representations in 2D rather than 3D due to limited high-quality 3D data availability [68]. Our object-centric rather than agent-centric visual aids approach effectively circumvents the limitations of heterogeneous embodied data while leveraging general visual datasets without robots, thus enabling robust generalization to novel scenarios and tasks.
Figure 3: Inspired by the process of human reasoning, FSD uses a spatial relationship graph as an anchor to derive a visual chain-of-thought reasoning process for visual trace generation.
3.2 Spatial Relationship-Focused Visual Chain-of-thought
To enable VLMs to generate spatial visual aids, a direct approach is supervised fine-tuning [39, 66, 23, 67] using these aids as a new action space or employing generative models [47, 59]. However, the heterogeneity and scarcity of embodied datasets limit this method. Directly aligning RGB images with coordinate points is challenging and prone to overfitting, hindering generalization. How can we stimulate VLMs’ spatial reasoning abilities to guide the generation rather than merely relying on extensive demonstrations? Inspired by human cognition (Fig. 3 (Top)), when executing tasks like "putting broccoli into a pot," humans first locate relevant objects, then plan movement paths based on relative positions while considering feasibility and obstacles. During this process, humans construct reasoning chains, repeatedly referencing object positions and establishing spatial relationships.
Based on these considerations, we introduce Spatial Relationship-Focused Visual Chain-of-thought (SrCoT). This approach guides VLMs to generate visual aids through structured reasoning based on spatial relationship graphs. SrCoT consists of two essential phases: 1 Description: We generate object-centric region captions establishing a task-relevant spatial relationship graph where nodes represent objects with their coordinates and edges denote relative relationships (above, below, left, right, etc.). 2 Reasoning: Using the spatial relationship graph as anchor points, we determine start and end coordinates through object references and free space reasoning, then iteratively derive intermediate points with explicit logical connections between steps. Thus, we prescribe a templated reasoning path for VLMs, enabling FSD to perform analogical reasoning in the spatial domain. While VLMs struggle to directly map future actions to image coordinates, our method leverages known object relationships as reference points for multi-hop analysis, simplifying the reasoning process. Fig. 3 (Bottom) demonstrates a complete reasoning sequence. This step-by-step SrCoT approach, though powerful, fundamentally depends on precise spatial understanding capabilities.
3.3 Weak-to-Strong Capability Dataset Construction Pipeline
The SrCoT mechanism demands enhanced capabilities from VLMs, including precise reference grounding, spatial understanding, and complex instruction-following capabilities (directly predicting future point trajectories) where mainstream VLMs [34, 30, 44, 8] show limitations. Consequently, we designed a weak-to-strong data construction pipeline to progressively develop these abilities.
FSD encompasses five hierarchical capability levels: 1 Region Grounding enables robots to focus on key objects in scenes. Although grounding capabilities have been broadly integrated into current VLMs [6, 64], understanding various small objects and complex scenes for embodied tasks is still limited; 2 Spatial Relationship understanding establishes prerequisite knowledge for spatial reasoning, forming the anchor points for SrCoT; 3 Spatial Reasoning builds upon these foundations to perform multi-hop analysis of object positions and relationships; and finally, 4 Spatial Affordance Generation and 5 Visual Trace Generation create actionable spatial guidance. Notably, SrCoT functions as a general visual-spatial reasoning mechanism applicable beyond visual traces to diverse spatial reasoning tasks. Through hierarchical spatial capability training, we enhance VLMs’ general spatial reasoning abilities, extending well beyond just embodied domains.
Figure 4: FSD screens data from large-scale embodied datasets, generates ground truth spatial relationship graph. We finally collected 300K data for 10+ embodiments with 5-level capabilities.
Automatic Dataset Construction: We leveraged extensive robot data from BridgeDataV2 [53], RT-X [40], and Droid [20] to construct FSD’s training dataset. Inspired by SpatialVLM [5] and SpatialRGPT [7], our automated data construction pipeline created 300K Supervised Fine-Tuning (SFT) data across numerous formats. After filtering demonstrations with unclear instructions, we used GPT4o [15] to nominate task-relevant objects while excluding out-of-range or overly complex items. We then built around objects using GroundedSam [45] for bounding boxes and segmentation masks (Level 1 dataset). For spatial relationship labels, we reconstructed 3D semantic scene graphs using Metric3Dv2 [13] for depth estimation, along with WildCamera [73] and PerspectiveFields [17] for camera parameters. This enabled 2D-3D mapping and spatial relationship graph construction. (Level 2 dataset). We only generate relative depth sorting data to infer positional relationships, so the accuracy requirement is not high. To improve data quality, in particular, we selected objects with a relative depth gap of 20% for subsequent generation. Afterward, we randomly sample spatial relationship graphs to construct spatial reasoning QA. We provide image captions, object coordinates, and relationships as a context for GPT4o to create complex QAs (Level 3 Dataset).
A core aspect of the FSD dataset is the visual aids generation. We employ a simple method with successful human demonstrations from embodied datasets and infer the process from the results. Spatial affordance represents the designated completion area for manipulation tasks. To create spatial affordance labels (Level 4 Dataset), we extract the manipulated object’s final position from the terminal frame, combine it with reference object positioning, calculate the precise affordance region, and re-render this information onto the initial frame. For visual trace generation (Level 5 Dataset), we employ a two-stage approach: first applying self-supervised keypoint extraction [14] to identify grasp points on manipulated objects, then utilizing Cotracker [19] to capture temporal dynamics from human demonstrations, subsequently projecting these trajectories onto the initial frame. Throughout this process, we employed strict rule-based filters and continually validated our approach against manually annotated test sets, iteratively refining our filtering criteria based on empirical feedback to ensure the resulting dataset met our quality requirements. The dataset presentation, data filtering process, and prompts used to generate the data are provided in Appendix A.
3.4 Self-Consistent Alignment for Spatial Understanding and Generation
High-quality SFT datasets enable VLMs to generate visual aids [66], yet these models struggle to understand the physical meaning of such annotations since coordinate spaces never appeared in pretraining data. The alignment between image coordinates and actual spatial positions presents a significant challenge. Therefore, we propose a self-consistency mechanism to further align FSD capabilities in spatial understanding and generation. We frame generation tasks inversely as understanding problems: if the forward task requires inferring visual trace $\boldsymbol{\tau}$ from an image $X_{v}$ and task instruction $X_{q}$, i.e. $(X_{v},X_{q})\rightarrow\boldsymbol{\tau}$, we construct the inverse task of predicting possible instructions given an image and visual traces $(X_{v},\boldsymbol{\tau})\rightarrow X_{q}$. This bidirectional approach helps the model comprehend spatial coordinates’ meanings and aligns coordinate space with image-text modalities, unifying visual aids as both understanding and generation signals while enhancing FSD spatial reasoning capabilities.
4 Training and Action Execution of FSD
Training: We follow the instruction tuning pipeline proposed by LLaVA-1.5 [33, 34]. As shown in Fig. 1, FSD’s architecture comprises an image encoder (CLIP-ViT-L-336px [11]), a linear projector, a language tokenizer, and a LLM (Vicuna-13B [71]). The image encoder processes images into tokens. These visual tokens are then projected into the same embedding space as language tokens through a two-layer linear projector. Only the projector and LLM weights are updated during fine-tuning while the vision encoder and tokenizer remain frozen. We built upon ASMv2 [55] as our foundation, which already incorporates basic relation conversation and reference grounding capabilities. The training process of FSD is divided into two stages: General Spatial Reasoning Enhancement: Using data from levels 1-3, we focus on improving the model’s embodied spatial reasoning capabilities. Following Yuan et al. [66] and Brohan et al. [3], we discovered that an appropriate data mixture is crucial for downstream performance. Joint training with mixed robotic and internet data ensures the model retains knowledge acquired during pre-training. Consequently, our instruction tuning utilizes a diverse 1.4M sample mixture including general visual question answering (VQA) data. This comprehensive training ensures FSD maintains robust general spatial knowledge while developing embodied capabilities. Visual Aids Generation and Understanding: Using data from levels 4-5 with the self-consistency mechanism, we specifically train visual aids generation and understanding abilities. FSD predicts a fixed set of 8 points for simplification when generating spatial visual traces. Additional training details and the summary of mixture datasets are provided in Appendix B.
Action Execution: FSD can reason from initial or intermediate task steps, freely selecting needed visual aids. When using bounding boxes, we sample the center as the target point; with affordance points, we directly sample one point. For visual trace execution, we first generate 2D visual traces $\boldsymbol{\tau}$ and obtain preliminary depth information from depth cameras. Following the pinhole camera model, we employ depth-based back-projection to map these to 3D space, yielding $\boldsymbol{\tau}^{3d}=\{\mathbf{x}_{t}^{3d}\mid t=1,2,...,T\}$. Next, based on the spatial position of the first point $\mathbf{x}_{1}$, we query GraspNet’s [9] grasp candidates $G$ to match the nearest grasp pose $G^{*}$. For relatively fixed scenes, we may also use predetermined grasp poses. Subsequently, we optimize the path trajectory using gradient descent-based interpolation, generating complete motion trajectories in SE(3) space, enabling the robotic arm to follow the 3D visual trajectory. When using only spatial affordance, we utilize CuRobo [49] as the motion planner to determine execution trajectories $T$ based on the target position $\mathcal{P}$. More details are provided in Appendix C. Unlike methods such as LLARVA [39] and EmbodiedCOT [67] which also utilize visual auxiliary aids, FSD transforms prediction tasks into reasoning tasks, better leveraging visual-spatial common knowledge without requiring scenario-specific fine-tuning.
5 Visual Aids Generation Benchmark
Few datasets exist for evaluating visual aid generation, Where2Place [66] provides 100 real-world images from homes and offices but is limited to direct and simple language instructions. And no benchmarks for trajectory prediction. To address this gap, we propose the Visual Aids Generation Benchmark (VABench). we manually annotated 300 problems from real-world and simulation datasets (OXE [40], BridgeData [53], Droid [20], Libero [31]), requiring models to infer visual aids given only natural language instructions similar to everyday human commands. For spatial affordance (VABench-Point), we measure the proportion of points falling within target regions. For models that only output bounding boxes, we sample uniformly within boxes. For visual trace (VABench-VisualTrace), we compute MAE and RMSE between ground truth $\boldsymbol{\tau}=\{\mathbf{x}_{t}\mid t=1,2,...,T\}$ and predictions $\hat{\boldsymbol{\tau}}=\{\hat{\mathbf{x}}_{t}\mid t=1,2,...,\hat{T}\}$: $\text{MAE}=\frac{1}{T}\sum_{t=1}^{T}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|$, $\text{RMSE}=\sqrt{\frac{1}{T}\sum_{t=1}^{T}\|\mathbf{x}_{t}-\hat{\mathbf{x}}_{t}\|^{2}}$. We interpolate when trajectory lengths differ and normalize coordinates to a 1000×1000 space for consistent evaluation. Since multiple valid solutions exist for each instruction, we simulated human evaluation standards by establishing detailed assessment criteria and added metrics with MLLM-based qualitative scoring (0-10) of visualized trajectories, named GPT Score. We provide detailed dataset information and evaluation procedures in Appendix D.
Figure 5: FSD directly generates visual aids based on task instructions for novel tasks and scenarios. 1st row: affordance bounding boxes; 2nd row: affordance points; 3rd and 4th rows: visual traces.
6 Experiments
We evaluated FSD across two dimensions: Seeing and Doing. For Seeing, we tested its general spatial reasoning and visual aids generation capabilities. For Doing, we conducted zero-shot manipulation experiments in both SimplerEnv [25] simulation and real-world xArm robotic platforms to assess its practical generalization performance.
6.1 Evaluation of Spatial Understanding and Reasoning Capabilities
Benchmarks and Baselines. General Spatial Reasoning Capabilities: We evaluated general spatial reasoning capabilities using five popular benchmarks: CVBench [51], BLINK [10], CRPE [55], SAT [44], and EmbSpatial-Bench [8]. These benchmarks encompass 15 subtasks measuring various spatial competencies including counting, spatial relationships, distance estimation and so on. We included two leading closed-source models: GPT-4o [15] and GPT-4V as performance reference. Subsequently, we conducted comparative analyses against other open-source spatial enhanced MLLMs, including LLaVA-1.5 [33], SAT-Dynamic [44], RoboPoint [66], and ASMv2 [55], all with 13B parameters. Object and Free Region Reference Capabilities: Following Yuan et al. [66], we assessed embodied spatial capabilities using the RoboRefIt [36] and Where2Place [66] benchmarks. We compared mainstream closed-source models and MLLM for enhancing spatial abilities (SpatialBot [4], SpaceLLaVA [5], and RoboBrain [16]). We used the proportion of predicted points within specified objects/regions as the accuracy metric. For models without point output support, we asked models to output bounding boxes of target objects/free space regions, then sampled evenly within these bounding boxes. Spatial Affordance and Visual Trace Capabilities: We utilized our VABench to evaluate the capabilities. We found very few models with this capability for Visual Trace prediction, so we trained an end-to-end prediction baseline model using a pre-trained DINOv2 [41] encoder coupled with a multi-layer transformer [52] architecture to predict visual trajectories, trained on the same visual trajectory data, named DINOv2 Predictor. We conducted this comparison to demonstrate the advantages of our reasoning-based FSD approach. More detailed descriptions of benchmarks and baselines are provided in Appendix E.
FSD exhibits superior general spatial reasoning capabilities. As shown in Table 1, FSD achieves a top ranking of 1.3 across 18 subtasks from 5 spatial benchmarks, surpassing other 13B VLMs and competing with GPT-4o. FSD demonstrates particular strengths in 3D depth perception (88.0%), 3D distance estimation (86.7%), and spatial relationship (78.3%). This confirms the effectiveness of our data mixing strategy and capability construction in enhancing spatial reasoning abilities. We believe stronger spatial foundational capabilities enable enhanced embodied perception.
FSD excels in object reference and free space localization. The results in Table 2 demonstrate FSD’s ability to accurately identify objects and free spaces from language instructions. For object reference (RoboRefIt), FSD achieves 56.7% accuracy, surpassing both GPT-4o (15.3%) and specialist models like RoboPoint (49.8%) by significant margins. On the more challenging free space reference task (Where2Place), FSD performs competitively with RoboPoint (45.8% vs. 46.0%) while substantially outperforming other models. This improvement stems from enhanced spatial understanding through our SrCoT mechanism. See Appendix F for more visualizations.
Table 1: Performance comparison on 5 spatial reasoning benchmarks. Bold and underlined values show best and second-best performance among open-source models.
CVBench CRPE SAT BLINK EmbSpatial Rank
Count 2DRel 3DDep 3DDis Avg. Exist. Subj. Pred. Obj. Avg. Val Real Count MV RelDepth SpRel Avg. Test
Closed-source models
GPT-4V 62.4 71.1 79.8 68.3 70.4 90.6 76.7 65.1 68.5 75.2 44.8 50.7 60.8 55.6 59.7 72.7 62.2 36.1 -
GPT-4o 65.9 85.5 87.8 78.2 79.4 93.3 81.9 71.8 73.6 80.2 49.4 57.5 49.2 60.2 74.2 69.2 63.2 49.1 -
Open-source models
LLaVA-1.5-13B 58.2 46.6 53.0 47.8 51.4 88.7 57.4 54.2 55.2 63.9 51.4 41.6 45.0 41.4 53.2 69.9 52.4 35.1 4.8
SAT-Dynamic-13B 61.5 89.7 80.7 73.0 76.2 87.5 60.6 57.6 65.2 67.7 87.7 54.9 35.8 44.4 73.4 66.4 55.0 51.3 2.8
RoboPoint-13B 56.5 77.2 81.5 57.7 68.2 93.2 66.3 62.4 70.9 73.2 53.3 46.6 48.3 44.4 62.1 65.7 55.1 51.4 2.8
ASMv2-13B 58.9 68.9 68.9 68.9 66.4 92.1 69.2 59.0 65.3 71.4 63.9 46.7 59.2 44.4 56.5 65.0 56.3 57.4 3.1
FSD-13B 62.4 86.5 88.0 86.7 80.9 94.0 75.2 65.1 70.4 76.2 73.2 63.3 60.0 46.6 70.2 78.3 63.8 63.3 1.3
Table 2: Performance comparison on object/free space reference benchmarks. The best results are highlighted in bold.
Benchmark GPT-4o SpaceLLaVA LLaVA-NeXT-34B SpatialBot-3B ASMv2-13B RoboBrain-7B RoboPoint-13B FSD-13B
RoboRefIt 15.3 21.3 19.9 23.6 48.4 10.1 49.8 56.7
Where2Place 29.1 11.8 15.0 15.0 22.0 16.6 46.0 45.8
Figure 6: Example of the FSD Reasoning Process FSD performs point-by-point reasoning and localization, ultimately generating a feasible visual trace.
Figure 7: Real-world robotic manipulation tasks performed by an xArm 6 robot (left) and performance comparison of FSD against baseline models GPT-4o and RoboPoint (right).
Table 3: Performance comparison on VABench. The best results are highlighted in bold.
(a) VABench-Point
Model Accuracy $\uparrow$
GPT4o 9.30
ASMv2 10.07
RoboPoint 19.09
FSD 61.82
   w/o SrCoT 26.21
   w/o Alignment 55.92
(b) VABench-VisualTrace
[…]
DINOv2 Predictor 128.32 117.49 2.17
FSD 78.26 63.44 5.49
   w/o SrCoT 99.53 80.06 4.00
   w/o Alignment 80.48 66.80 5.45
Table 4: SimplerEnv Evaluation on WidowX Robot The results of baselines are derived from [43]. ZS: zero-shot, FT: fine-tuning using BridgeData. Each task is tested 24 episodes.
Model Spoon→Towel Carrot→Plate Green→Yellow Eggplant→Basket Avg.
Grasp Succ. Grasp Succ. Grasp Succ. Grasp Succ.
RT-1-X [40] 16.7% 0% 20.8% 4.2% 8.3% 0% 0.0% 0% 1.1%
Octo-S [50] 77.8% 47.2% 27.8% 9.7% 40.3% 4.2% 87.5% 56.9% 30.0%
OpenVLA [21] 4.1% 0% 33.3% 0% 12.5% 0% 8.3% 4.1% 1.0%
RoboVLM (ZS) [24] 37.5% 20.8% 33.3% 25.0% 8.3% 8.3% 0.0% 0% 13.5%
RoboVLM (FT) [24] 54.2% 29.2% 25.0% 25.0% 45.8% 12.5% 58.3% 58.3% 31.3%
SpatialVLA (ZS) [43] 25.0% 20.8% 41.7% 20.8% 58.3% 25.0% 79.2% 70.8% 34.4%
SpatialVLA (FT) [43] 20.8% 16.7% 29.2% 25.0% 62.5% 29.2% 100% 100% 42.7%
FSD 58.3% 41.6% 58.3% 50.0% 91.6% 33.3% 37.5% 37.5% 54.1%
FSD demonstrates breakthrough performance in visual aid generation. As shown in Table Table 3(b), FSD significantly outperforms all baseline methods in generating precise spatial affordances and visual traces. For VABench-Point, FSD achieves 61.82% accuracy—three times higher than RoboPoint (19.09%), while for VABench-VisualTrace, FSD demonstrates significantly lower error rates and better GPT Score compared to the DINOv2 Predictor. Ablation studies confirm the critical contributions of both SrCoT and self-consistency alignment, validating our reasoning-based approach where step-by-step spatial analysis enables more accurate predictions than purely data-driven methods. We provide detailed comparisons with RoboBrain [16] in Appendix I.
Visualization of visual aids generated by FSD We present examples of generation in Fig. 5. FSD is able to adapt to diverse scenes, perspectives, and tasks, and the three types of visual traces generated can well match the task instructions. We also presented the reasoning process of FSD in Fig. 6, where the CoT provides clear guidance for the results. More visualizations are in Appendix F.
6.2 Evaluate the Decision Capability
SimplerEnv Simulation SimplerEnv [25] is a simulation environment specifically designed for evaluating real-world robotic manipulation. It provides a standardized platform, emphasizing reproducibility with real-world scenarios. We utilized FSD to generate visual traces and perform zero-shot deployment on the WidowX robotic arm. We compared its performance against state-of-the-art general manipulation VLA, including Octo [50], OpenVLA [21], RoboVLM [24] and SpatialVLA [43]. The results are presented in Table 4. FSD surpasses the baseline in most tasks without FT with a 54.1% success rate, demonstrating the strong generalization ability of FSD. Notably, FSD achieves the best average performance, even outperforming VLA models that even outperforming specially fine-tuned VLA models. This superior generalization capability demonstrates that visual aids representations provide better zero-shot transfer compared to direct action prediction approaches, bridging the gap between perception and control more effectively.
Real World Robot Evaluation As shown in Fig. 7, we conducted zero-shot tests with FSD on an xArm 6 robot across 8 tabletop manipulation tasks. The setup included an Intel RealSense L515 LiDAR camera. To test the capabilities of different visual aids, we used visual trace for the sponge and folding tasks, while affordance points were used for other tasks. We compared against RoboPoint and GPT-4o baselines, which were limited to predicting only start/end points rather than full trajectories. Robopoint often makes incorrect predictions in tasks involving spatial understanding. Under zero-shot conditions, FSD achieved 72% success rate, outperforming the strongest baseline by more than 30%. Notably, FSD successfully completed complex tasks with visual trace generation, e.g. cloth folding, which was beyond baseline capabilities. Full results are presented in Fig. 7 and we refer to Appendix G for detailed setup and more visualizations.
7 Conclusion
In this paper, we introduced FSD (From Seeing to Doing), bridging visual reasoning and robotic manipulation through intermediate spatial representations. Our approach addresses the challenges of data scarcity and embodiment heterogeneity through three key innovations: Spatial Relationship-Focused Visual Chain-of-Thought for multi-step reasoning; A hierarchical weak-to-strong data pipeline; and A self-consistency mechanism aligning spatial coordinates with visual signals. Experiments demonstrate FSD’s superior performance across multiple spatial reasoning visual aid generation benchmarks. In zero-shot robotic deployment, FSD achieved 72% success rate across diverse tasks, outperforming baselines by 30%. Limitations include reliance on 2D rather than 3D trajectory generation and constraints from available training data quality. Future works are detailed in Appendix J.